1.) If the Kb of a weak base is 3.4 x 10^-6, what is the pH of a 0.11M solution of this base?

2.) A certain weak base has a Kb of 7.00 x 10^-7. What concentration of this base will produce a pH of 10.16?
Answer #1

Let weak base be BOH

               BOH (aq) <---> B+ (aq) + OH- (aq)

at equilibrium [BOH] = 0.11-X , [B+] =[OH-] = X

Kb = [B+] [OH-] /[BOH]

3.4 x 10^-6 = X^2 / ( 0.11-X)

X = [OH-] = 0.0006 ,   pOH =-log [OH-] = -log ( 0.0006) = 3.22

pH = 14-3.22 = 10.8

2)   here pH = 10.16 , pOH = 14-10.16 = 3.84 , [OH-] = 10^-pOH = 10^- 3.84 = 0.0001445 M = [B+]

we Kb = [B+] [OH-] /[BOH[

7 x 10^-7 = (0.0001445) ( 0.0001445) / [BOH]

[BOH] = 0.03

[BOh] initial = [BOH] at equilibrium + [BOH dissociated to OH-]

      = 0.03 + 0.0001445 M

    = 0.03 M    ( rounded value)

Thus concentration of base required = 0.03M